OpenStudy (anonymous):

What is 15.5/100 in simplest form

OpenStudy (anonymous):

You can view the .5 = \[\frac{ 1 }{ 2}\]

OpenStudy (anonymous):

so you will now end up with \[\frac{ 15+\frac{ 1 }{ 2 } }{ 100 }\]

OpenStudy (anonymous):

You should be able to make the 15 and 1/2 into 1 fraction with a common denominator and after you do that you would just have to divide that by 100

OpenStudy (anonymous):

If you are not able to do it, post back with your work and someone will help you through it. Good luck

OpenStudy (anonymous):

15.5/100 in simplest form is 31/200

OpenStudy (anonymous):

That is true and I hope that MathNinja197 arrived at the answer before it was given so that MathNinja197 was not robbed of the understanding